Circular No. 8024

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hh IN NGC 6946 P. Meikle, Imperial College, London; S. Mattila, Stockholm Obser-

vatory; and S. Smartt, Institute of Astronomy, Cambridge, report that infrared observations of the type-II supernova 2002hh obtained by E. Mac-Donald, L. Clewley, and G. Dalton at the William Herschel Telescope (+ infrared camera INGRID) on Nov. 18.86 UT yield J = 12.30 ± 0.03, K_{s} = 11.07 ± 0.02. Adopting a phase of 21 ± 2 days post-explosion, comparison of the J − K_{s} color with the infrared-template light curves of Mattila and Meikle (2001, MNRAS 324, 325) indicates that the supernova is highly reddened, with E(J − K_{s}) = 1.0, confirming the result of Filippenko et al. (IAUC 8007). Using the reddening law of Cardelli et al. (1989, Ap.J. 345, 245), this yields A(V) ∼ 6.1. Subtracting the Galactic extinction (Schlegel et al. 1998, Ap.J. 500, 525) yields a host-galaxy extinction A(V) ∼ 5.0. Adopting a distance of 5.9 Mpc (H_{0} = 70; Tully 1988, Nearby Galaxies Catalog) yields de-reddened absolute magnitudes of M(J) = −18.3 and M(K_{s}) = −18.5, which are close to the values given by the templates of Mattila and Meikle (2001).

D. Pooley and W. H. G. Lewin, Massachusetts Institute of Technology, on behalf of a larger collaboration, report the detection of x-ray emission from SN 2002hh with the Chandra x-ray observatory: “A 30000-s ACIS-S3 observation on Nov. 25.36 UT revealed a source at α = 20^{h}34^{m}44^{s}.24, δ = +60^{o}07′19′′.3 (equinox 2000.0; ± 0′′.5 in each coordinate), in close agreement with the reported optical and radio positions. A previous Chandra observation of NGC 6946 on 2001 Sept. 7 (60000 s with ACIS-S3) shows that no x-ray source was present at this location to a 3σ limiting flux of roughly 7 × 10^{−16} erg s^{−1} cm^{−2} in the band 0.4–8 keV, corresponding to a limiting luminosity of 3 × 10^{36} erg/s for a distance of 6.3 Mpc (Tully 1988, Nearby Galaxies Catalog, using H_{0} = 65 km/s/Mpc). Preliminary spectral fits to the recent Chandra data (with both MEKAL and power-law models) indicate a rather hard, highly absorbed spectrum (as expected from the high reddening reported on IAUC 8007), with a fitted column density of N_{H} = 10^{22} cm^{−2} (compared to N_{H} = 2–3 × 10^{21} cm^{−2} reported for just NGC 6946 by Burstein and Heiles 1984, Ap.J. Suppl. 54, 33). The powerlaw photon index is found to be 0.6, but the MEKAL temperature could not be accurately measured because the data prefer a value of kT above 80 keV, which is the model limit. The unabsorbed 0.4–8-keV luminosity calculated from both models agrees well and is 5 × 10^{38} erg/s. This low x-ray luminosity supports the indication on IAUC 8018 that there is little circumstellar interaction taking place.”

2002 November 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green